36-726: Statistical Practice On working with your investigator

In your stat practice project, you are taking on a new role that many of you are unfamiliar with: you are working as a kind of collaborator with the faculty investigator who brought his or her project to the class. This is a professional relationship: you must show appropriate courtesy, respect and etiquette in all aspects of this relationship. That means both social courtesy and the intellectual respect of both working hard to understand what your collaborator (the faculty investigator) wants and needs, and communicating clearly and courteously your best advice, feedback and analysis for the project.

Please note the following guidelines for working with your faculty collaborators:

1. Plan to meet with them about once a month. Some projects may naturally lead to more frequent meetings but you shouldn’t meet any less than once a month. Be sensitive to the schedule of your collaborators, in frequency of meetings (and all things).

2. Schedule meetings well in advance. ALWAYS be up front, courteous and flexible about scheduling meetings, and ALWAYS be on time for meetings.

3. Be professional, polite and courteous at all times in meetings (with your teammates, and with your scientific collaborator). Your collaborators are accustomed to and are expecting professional collaborators. This doesn’t mean you should be meek. It means you should be socially graceful. For example, it is possible (and necessary, across scientific and academic cultures) to be graceful while digging to get to the bottom of intellectual, organizational or other questions.

4. ALWAYS have paper and pencil (or laptop if you are good at using one) out during meetings, and ALWAYS be taking copious notes. Scientists and other professionals love to talk about their work, so you won’t have any trouble getting them to talk. Afterwards, study and organize your notes, so you know what to follow up on (or not) in future meetings, email, phone calls, etc. The times your collaborators will get impatient if they have to give the same lecture twice or three times. On the other hand, engaging an old issue with new or deeper questions or in new directions is an activity your collaborators will enjoy.

5. Find out how your collaborators like to communicate between meetings. Some are more tuned to email, some are more tuned to the phone. Be sensitive to how often your collaborators want to be contacted between meetings. Some people like questions saved up in batches that only come occasionally, others like to address each little question as soon as you think of it. People who like one style typically don’t like the other. Be flexible yourself and adapt to your collaborator’s style.

Always feel free to come talk to me about about any aspect of the project. We will set up regular meetings for you and me in the spring, but don’t wait for our regular meeting if you are concerned or uncertain about anything (social, scientific, statistical, etc.). Drop by my office or send email. If I don’t know the answer I’ll help you find it.

Brian Junker 132E Baker Hall 268-8874 brian@stat.cmu.edu

Summary of Some Important Points on Consulting

• Welcoming Setting / Body Language / Social Interaction

– Setting the agenda, and allowing for flexibility

– Accomodating client’s communication style / thinking style

• Gathering Information

– Open-ended, natural language questions, vs. closed and/or technical questions

– When technical terms are unavoidabe, explain them and make sure they are understood in the client’s terms

– Paraphrasing/summarizing client’s input – make sure client agrees that he/she has communicated!

• Client as Collaborator

– Presenting or surfacing problems for client to solve, without providing support for solving them

– Leaving client uninvolved in decision making.

– Anticipating and solving problems for client, and/or providing support tools to help solve problem, e.g.: ∗ Institutional resources (when this makes sense)

∗ Computational tools consumable by client

∗ Summaries of information consumable by client – Maintaining client involvement in intellectual enterprise and decision making.

• Delivery of Statistical Information

– help client understand the information you are communicating, or the process or product that you are

proposing or reporting on ∗ Explain terms

∗ Use visual aids

∗ Use computer aids

∗ Use other aids

∗ Use client’s background

∗ Etc.

– Deliver information formatted so that audiences with different needs can easily consume it – negotiating about needs helps!

• Negotiating

– Clarifying and defining the project: tasks, goals, schedule, budget, possible uses of data (by both sides!)

– Clarifying the responsibilities and commitments of both sides

– Clarifying the expectations of both sides – e.g., presenting examples of what each side hopes to get from the relationship

– This is ongoing throughout the project, not just at one meeting!

• Closing

– Summarize tasks/responsibilities/expectations of client and consultant.

• Breakdowns

– Occur because expectations are not verbalized, agreed upon, or simply not met.

– They happen in every consulting project

– Learning how to recognize and resolve them is an important skill!

– If you maintain lines for negotiation and frank discussion of problems as they arise, big breakdowns at the end are less likely.